# Optimizing Power Grid Topologies with Reinforcement Learning: A Survey of Methods and Challenges

### v1

@@ 2 Overview of Challenges and Solutions
This chapter provides a comprehensive overview of the L2RPN competitions, highlighting the most significant contributions to each challenge. [10] has been used to help create this overview. Several solutions proposed outside the L2RPN competitions are also discussed in the last part of this section. For a broader understanding and detailed description of the challenges, we direct readers to [6] and the articles mentioned in the specific sections. For many of the proposed solutions, we introduce acronyms for easier reference and recognition throughout the paper. These acronyms follow the format AAA-YYYY, where AAA represents a few descriptive letters and YYYY indicates the year. We provide an overview of all the acronyms used and references in Table 33 in Appendix A.
@@ 2.1 Competition and solutions L2RPN 2019 IJCNN - Sandbox
The first challenge [1], introduced in 2019, was based on the IEEE 14-bus system. The platform used for this challenge was PyPowNet [4], which has been replaced with Grid2Op in subsequent challenges. In this challenge, the goal was to design an agent that can operate a power grid in real-time over a one-month period using only topology changes. That is, only bus-splitting and line-switching actions were used.
This competition had two winning solutions [11] (A3C-2019), [12] (DDQN-2019).
A3C-2019 is the only solution that does not use the Grid2Op simulation function and is, therefore, purely RL-based. It uses an asynchronous-advantage-actor-critic (A3C) algorithm [13] with a curriculum learning strategy [14], where the idea is that the neural networks learn more effectively when they are first trained on a simpler task. Furthermore, the authors reduce the action space by eliminating all busbar symmetry actions within a substation, since these result in an identical situation and are therefore redundant. This is something that, in subsequent solutions, is done by default and is included when collecting all substation topology actions in the Grid2Op package.
The solution DDQN-2019 uses imitation learning (IL), [15], combined with a Dueling DQN algorithm, [16], that is trained using importance sampling or prioritized experience replay, [17]. They introduce a warning flag based on the loading of a line level. This warning flag can be seen as an activation threshold that determines when the agent should interfere with the network to avoid further escalation. When the load on all lines is below the activation threshold, the agent is not activated, but a simple do-nothing action is given to the environment. In all subsequent solutions, an activation threshold is used. In addition, guided exploration is used to speed up training. The final solution simulates the top $N$ actions given by the agent and takes the best action.
@@ 2.3 L2RPN 2020 NeurIPS - In a sustainable world
[30] (Binbinchen-2020) won the second-place for the robustness challenge. The authors propose a Teacher-Tutor-Junior-Senior framework. Here, the Teacher first defines a smaller action space of 208 actions by applying a greedy agent to the entire action space and saves all used greedy actions in a reduced action space (RAS). The RAS is used by the Tutor, the tutor applies actions, again greedily, and the Junior updates a neural network using IL. Ultimately, the Senior employs this neural network to initialize the actor network for the Proximal Policy Optimization (PPO) algorithm [31].
@@ 2.4 L2RPN 2021 ICAPS - Run a power network with trust
The third place in the competition was secured by [37]. This solution served as a precursor to their winning approach for the 2022 WCCI competition AlphaZero-2022 ([38]) also using the AlphaZero approach with Monte Carlo Tree Search (MCTS) sampling [39] as the core RL algorithm. Notably, in this solution, they preserved the original action space to avoid losing potential beneficial actions. A benefit of their solution is that MCTS already requires looking into the future and, therefore, helps identify critical situations early. This capability is then used for the alarm strategy.
@@ 2.5 L2RPN 2022 WCCI and 2023 TU Delft - Energies of the Future and carbon neutrality
In the challenge of 2022 [5], the energy mix is changed to obtain more realistic future scenarios to reach carbon neutrality in 2050; 40% of the energy mix consists of renewables; wind and solar power and less than 3% is fossil fuels. The use of renewable energy generators is privileged and the use of fossil fuels is penalized to steer agents. The results of the competition can be found in the blog post of [40].
As mentioned in the previous section, [37] improved their solution submitted to the L2RPN 2021 ICAPS competition. Unlike their previous solution, the action space is now narrowed down to 2000 actions using a brute-force selection method. In addition to the learned topology agent, they introduce a redispatching controller that uses Cross-Entropy (CE) optimization [41]. Topology actions are preferred over redispatch actions, but when both cannot solve the congestion, they rerun the redispatch optimization on the top five proposed topology action candidates. With the improved solution AlphaZero-2022, [38] won the 2022 WCCI competition. The authors also wrote a follow-up paper on how to integrate such an RL-based agent into existing grid operation workflows, [42].
@@ 2.6 Solutions outside of competitions
In the paper of [49] (CEM-2021), the Cross-Entropy Method [41] is applied to a modified version of the 14-bus power grid, case_14_realistic in Grid2Op. With this simple RL approach and a smart selection of episodes, the authors were able to train an agent to successfully operate $96.5\%$ of the scenarios consisting of one week. The research considers bus-splitting actions only and proposes a reduction of the action space by excluding actions where only one element is connected to a busbar and actions where none of the connected elements is a line (these are infeasible actions and are currently excluded by default when collecting topology actions in Grid2Op). Furthermore, they show how the number of actions per substation can be computed. An interesting part of this research is the evaluation of the agent, which shows that the agent only needs very few different topologies to manage the grid. Their agent manages the grid with topology actions at substations 1, 3, and 8.
@@
[54] introduce a hierarchical multiobjective Markov decision problem (HMO-MDP) to address more fairness between the power plants. This is done by introducing a suitable reward function and is tested using A2C and PPO.
@@
A recent study on IL for PNC, [62], introduces a soft label IL approach (SoftIL-2025). Instead of directly learning from greedy-based action selection, as in previous methods, all actions are labeled based on the line loads during the simulations, providing a richer learning signal for the agent. They evaluate this method using FCNN and GNN architectures on the 118-bus case (WCCI 2022 L2RPN). The soft-labeled GNN, in particular, achieves strong performance, outperforming current state-of-the-art RL agents.
@@ 3 Design choices for RL in Power Grid Control
Solution Acronym Competition Size Grid With Opponent Rank Algorithm used Imitiation learning Prioritized replay Simulation used Activation threshold Reconnect lines Revert topology A3C-2019 IJCNN 2019 14 1/2 A3C 0.6  [^3] Based on code. DDQN-2019 IJCNN 2019 14 1/2 DDQN 0.885 SMAAC-2021 WCCI 2020 14, 36 1 SAC 0.9 A3C-2020 WCCI 2020 36 3 A3C 0.8 SAS-2021 NeurIPS 2020 36, 118 1 ES 1.0 444 This activation threshold is compared with the $\rho_{\max}(t+1)$ after a simulated DN-action. So unlike other solutions where the current $\rho_{\max}(t)$ is checked. Binbinchen-2020 NeurIPS 2020 36 2 PPO 1.0 (Teacher) 0.925 (Tutor) 0.9 (Senior) 0.999 (final agent) D3QN-2020 NeurIPS 2020 36, 118 3 D3QN 1.0 3 AlphaZero-2022 WCCI 2022 118 1 MCTS + CE 0.98  [^5] Although not explicitly documented, this value was used in their ICAPS 2021 solution, suggesting it was applied again here. BruteForce-2022 WCCI 2022 118 2 Greedy + OptimCVXPY 0.95 HRI-EU-2022 WCCI 2022 118 3 Greedy 1.0 4 LJNAgent-2024 WCCI 2023 118 1 PPO + OptimCVXPY $[0.6;1.99]$ (Teachers) 0.99 (final agent) Artelys-2024 WCCI 2023 118 2 PPO + OptimCVXPY 0.98 CEM-2021 - 14 CEM 0.95 D3QN-2022 - 14 D3QN 0.95 PowRL-2022 - 36 PPO 0.95 4  [^6] The specific activation threshold is not documented, but given that the reward function uses a safety threshold of 0.95, likely, this value was also applied as the activation threshold. Curriculum-2023 - 36 PPO 0.925 (Teacher) 0.9 (Tutor) 0.9 (Senior) 0.95 (final agent) 3 HRL-2023 - 14 PPO / SAC 0.95 MARL-2023 - 5 MA-PPO / MA-SACD 0.9 HUGO-2024 - 118 PPO 0.85 (TopoAgent) 0.95 (final Senior) IL-2024 - 14 IL 0.97 HMO-2024 - 14, 118 A2C/PPO N.A. [^7] Not Available. It is either not documented or not used. N.A. N.A. BDQN-2024 - 14, 36 ApeX-DQN N.A. N.A. N.A. Zonal-2024 - 118 PPO 0.9 N.A. N.A. GNNIL-2025 - 14 IL 0.97 CCMA-2025 - 5, 14 PPO 0.95 SoftIL-2025 - 118 IL 0.9
@@ 3.1 General RL Techniques
3.2 Algorithm
As shown in Table 1, the most frequently used RL algorithm for PNC is PPO, featured in 11 of the 26 solutions reviewed in this overview, significantly more often than any other algorithm. Studies comparing PPO with other algorithms, such as SAC and A2C, consistently show that PPO performs better in PNC [53, 22, 54]. Note, however, that the agent’s performance depends not only on the choice of algorithm but also on other implementation details, which are discussed in the following paragraphs. Additionally, various versions of DQN have been applied to the PNC problem. In [50] (D3QN-2022), the authors demonstrate that the prioritized D3QN, an advanced variant of DQN, significantly speeds up the training.
3.3 Action space
As outlined in Section 1, the actions available to operate the power network include bus splitting, line switching, generator re-dispatch or curtailment and, optionally, storage changes (introduced in [5]). Topological actions – bus splitting and line switching – serve as a low-cost option to alleviate thermal overloads on the network [6]. However, the large, non-linear, and combinatorial nature of the action space makes optimal control of the grid topology beyond current state-of-the-art capabilities [1, 7], leaving these topological actions as a significant, yet underutilized, source of operational flexibility in power networks.
The inability of existing methods to navigate such a vast action space has motivated the L2RPN competitions [1, 6, 7], where the primary goal is to use RL, to explore the full range of switching options to control power flows. As a result, topological actions are the main focus of agents used in L2RPN. Footnote 10 provides an overview of the number of feasible bus-splitting and line-switching actions per case. It shows how the number of bus-splitting actions increases drastically when going from case 14 to case 36.
@@ 3.3 Action space
also presented by [52]. Note that this computation also includes the actions of substations where only one configuration is possible due to connectivity constraints. Therefore, the result of this computation is slightly higher compared to what is presented in Footnote 10, which excludes these actions.
@@
Table 3: Action space reductions for case 14, the PyPowNet, realistic, and sandbox version. The first two solutions use case 14 in PyPowNet, which is not mentioned in Footnote 10.
In Tables 4 and 5 we omitted the “Original Size Action Space” column, as these solutions primarily focus on bus-splitting actions, $\mathcal{A}_{bus}$, for which the action space sizes are already detailed in Footnote 10. As can be seen in Table 4, the exceptions to this are A3C-2020 and D3QN-2020.
@@
Solution Acronym Actions for RL agent Reduced Size Action Space % of actions included Reduction method Actions for rule-based heuristics SMAAC-2021 Bus splitting 66674 99.8% Mask reduction with $N=5$. Line switching. A3C-2020 Topology actions 300 0.4% 200 based on experience, 100 randomly selected. In addition, the action space contains 295 line-switching actions. Line switching. SAS-2021 Bus splitting 732  [^14] Two action spaces are used, $\mathcal{A}_{normal}$, which is used by default and $\mathcal{A}_{exception}$, which is used only when specific lines are disconnected. Here, $|\mathcal{A}_{normal}|=500$ and $|\mathcal{A}_{exception}|=232$. 1.0% N.A. Line switching and redispatching. Binbinchen-2020 Bus splitting 208  [^15] This action space includes change bus actions. This is different from set bus actions which most solutions use. See https://grid2op.readthedocs.io/en/latest/. 0.3% Greedy reduction. Line switching when line is disconnected. D3QN-2020 Topology actions or redispatch actions 786 15  [^16] In addition, the action space contains 58 line switching, 40 redispatching and one explicit do-nothing action. 1.2% Expert reduction. Line switching when line is disconnected. PowRL-2022 Bus splitting 240 0.4% Greedy reduction. Line switching. Curriculum-2023 Bus splitting 508 15 0.8% Greedy reduction. Line switching. BDQN-2024 N.A., but only bus splitting actions seem to be used 1080 1.6% Stable Topo reduction. N.A.
@@
Solution Acronym Actions for RL agent Reduced size action space % of actions included Reduction method Actions for rule-based heuristics SAS-2021 Bus splitting 1000 1.0% N.A. Line switching. D3QN-2020 Topology actions 978 1.4% Expert reduction15. In addition, the action space contains 185 line switching and one do-nothing action. Line switching. AlphaZero-2022 Bus splitting 2000 2.7% Greedy reduction. Line switching and redispatching. BruteForce-2022 314 0.4% Greedy reduction. Bus splitting, line switching and redispatching. HRI-EU-2022 N.A. N.A. Randomly draw 1000 line switch combined with redispatch actions each time the grid is in danger. Line switching combined with redispatch. LJNAgent-2024 Bus splitting 1516 2.1% Greedy reduction. Line switching and redispatching. Artelys-2024 Bus splitting 500 0.7% Greedy reduction. Line switching and redispatching. HUGO-2024 Bus splitting 2030 2.8% Apply the action space of AlphaZero-2022 with an additional 30 expert actions selected by RTE (Expert reduction). Stable Topo Reduction for the target topologies (TTs). Line switching. SoftIL-2025 Bus splitting 2030 2.8% Apply the action space of HUGO-2024. Line switching.

### v2

@@ 2 Overview of Challenges and Solutions
This chapter provides a comprehensive overview of the L2RPN competitions, highlighting the most significant contributions to each challenge. [37] has been used to help create this overview. Several solutions proposed outside the L2RPN competitions are also discussed in the last part of this section. For a broader understanding and detailed description of the challenges, we direct readers to [11] and the articles mentioned in the specific sections. For many of the proposed solutions, we introduce acronyms for easier reference and recognition throughout the paper. These acronyms follow the format AAA-YYYY, where AAA represents a few descriptive letters and YYYY indicates the year. We provide an overview of all used acronyms and the paper references in Table 33 in Appendix A.
@@ 2.1 Competition and solutions L2RPN 2019 IJCNN - Sandbox
The first challenge, introduced in 2019 and described in [1], was based on the IEEE 14-bus system. The platform used for this challenge was PyPowNet ([32]), which has been replaced with Grid2Op in subsequent challenges. In this challenge, the goal was to design an agent that can operate a power grid in real-time over a one-month period using only topology changes. That is, only bus-splitting and line-switching actions were used.
This competition had two winning solutions by [38] (A3C-2019), and [39] (DDQN-2019).
A3C-2019 is the only solution that does not use the Grid2Op simulation function and is, therefore, purely RL-based. It uses an asynchronous-advantage-actor-critic (A3C) algorithm by [40] with a curriculum learning strategy, see [41], where the idea is that the neural networks learn more effectively when they are first trained on a simpler task. Furthermore, the authors reduce the action space by eliminating all busbar symmetry actions within a substation, since these result in an identical situation and are therefore redundant. This is something that, in subsequent solutions, is done by default and is included when collecting all substation topology actions in the Grid2Op package.
The solution DDQN-2019 uses imitation learning (IL), see [42], combined with a Dueling DQN algorithm, [4], that is trained using importance sampling or prioritized experience replay, [43]. They introduce a warning flag based on the loading of a line level. This warning flag can be seen as an activation threshold that determines when the agent should interfere with the network to avoid further escalation. When the load on all lines is below the activation threshold, the agent is not activated, but a simple do-nothing action is given to the environment. In all subsequent solutions, an activation threshold is used. In addition, guided exploration is used to speed up training. The final solution simulates the top $N$ actions given by the agent and takes the best action.
@@ 2.3 L2RPN 2020 NeurIPS - In a sustainable world
[55] (Binbinchen-2020) won the second-place for the robustness challenge. The authors propose a Teacher-Tutor-Junior-Senior framework. Here, the Teacher first defines a smaller action space of 208 actions by applying a greedy agent to the entire action space and saves all used greedy actions in a reduced action space (RAS). The RAS is used by the Tutor, the tutor applies actions, again greedily, and the Junior updates a neural network using IL. Ultimately, the Senior employs this neural network to initialize the actor network for the Proximal Policy Optimization (PPO) algorithm by [6].
@@ 2.4 L2RPN 2021 ICAPS - Run a power network with trust
The third place in the competition was secured by [60]. This solution served as a precursor to their winning approach for the 2022 WCCI competition AlphaZero-2022 ([61]) also using the AlphaZero approach with Monte Carlo Tree Search (MCTS) sampling by [7], as the core RL algorithm. Notably, in this solution, they preserved the original action space to avoid losing potential beneficial actions. A benefit of their solution is that the MCTS tree already requires looking into the future and, therefore, helps identify critical situations early. This capability is then used for the alarm strategy.
@@ 2.5 L2RPN 2022 WCCI and 2023 TU Delft - Energies of the Future and carbon neutrality
In the challenge of 2022 introduced in [33], the energy mix is changed to obtain more realistic future scenarios to reach carbon neutrality in 2050; 40% of the energy mix consists of renewables; wind and solar power and less than 3% is fossil fuels. The use of renewable energy generators is privileged and the use of fossil fuels is penalized to steer agents. The results of the competition can be found in the blog post of [62].
As mentioned in the previous section, [60] improved their solution submitted to the L2RPN 2021 ICAPS competition. Unlike their previous solution, the action space is now narrowed down to 2000 actions using a brute-force selection method. In addition to the learned topology agent, they introduce a redispatching controller that uses Cross-Entropy (CE) optimization by [30]. Topology actions are preferred over redispatch actions, but when both cannot solve the congestion, they rerun the redispatch optimization on the top five proposed topology action candidates. With the improved solution AlphaZero-2022, [61] won the 2022 WCCI competition. The authors also wrote a follow-up paper on how to integrate such an RL-based agent into existing grid operation workflows, [63].
@@ 2.6 Solutions outside of competitions
In the paper of [70] (CEM-2021), the Cross-Entropy Method by [30] is applied to a modified version of the 14-bus power grid, case_14_realistic in Grid2Op. With this simple RL approach and a smart selection of episodes, the authors were able to train an agent to successfully operate $96.5\%$ of the scenarios consisting of one week. The research considers bus-splitting actions only and proposes a reduction of the action space by excluding actions where only one element is connected to a busbar and actions where none of the connected elements is a line (these are infeasible actions and are currently excluded by default when collecting topology actions in Grid2Op). Furthermore, they show how the number of actions per substation can be computed. An interesting part of this research is the evaluation of the agent, which shows that the agent only needs very few different topologies to manage the grid. Their agent manages the grid with topology actions at substations 1, 3, and 8.
@@
[75] introduce a hierarchical multiobjective Markov decision problem (HMO-MDP) to address more fairness between the power plants. This is done by introducing a suitable reward function and is tested using (Advantage Actor-Critic) A2C and PPO.
@@
A recent study on IL for PNC, [83], introduces a soft-label IL approach (SoftIL-2025). Instead of directly learning from greedy-based action selection, as in previous methods, all actions are labeled based on the line loads during the simulations, providing a richer learning signal for the agent. They evaluate this method using FCNN and GNN architectures on the 118-bus case (WCCI 2022 L2RPN). The soft-labeled GNN, in particular, achieves strong performance, outperforming current state-of-the-art RL agents.
@@ 3 Design choices for RL in Power Grid Control
Solution Acronym Competition Size Grid With Opponent Rank Algorithm used Imitiation learning Prioritized replay Simulation used Activation threshold Reconnect lines Revert topology A3C-2019 IJCNN 2019 14 1/2 A3C 0.6  [^4] Based on code. DDQN-2019 IJCNN 2019 14 1/2 DDQN 0.885 SMAAC-2021 WCCI 2020 14, 36 1 SAC 0.9 A3C-2020 WCCI 2020 36 3 A3C 0.8 SAS-2021 NeurIPS 2020 36, 118 1 ES 1.0 555 This activation threshold is compared with the $\rho_{\max}(t+1)$ after a simulated DN-action. So unlike other solutions where the current $\rho_{\max}(t)$ is checked. Binbinchen-2020 NeurIPS 2020 36 2 PPO 1.0 (Teacher) 0.925 (Tutor) 0.9 (Senior) 0.999 (final agent) D3QN-2020 NeurIPS 2020 36, 118 3 D3QN 1.0 4 AlphaZero-2022 WCCI 2022 118 1 MCTS + CE 0.98  [^6] Although not explicitly documented, this value was used in their ICAPS 2021 solution, suggesting it was applied again here. BruteForce-2022 WCCI 2022 118 2 Greedy + OptimCVXPY 0.95 HRI-EU-2022 WCCI 2022 118 3 Greedy 1.0 5 LJNAgent-2024 WCCI 2023 118 1 PPO + OptimCVXPY $[0.6;1.99]$ (Teachers) 0.99 (final agent) Artelys-2024 WCCI 2023 118 2 PPO + OptimCVXPY 0.98 CEM-2021 - 14 CEM 0.95 D3QN-2022 - 14 D3QN 0.95 PowRL-2022 - 36 PPO 0.95 5  [^7] The specific activation threshold is not documented, but given that the reward function uses a safety threshold of 0.95, likely, this value was also applied as the activation threshold. Curriculum-2023 - 36 PPO 0.925 (Teacher) 0.9 (Tutor) 0.9 (Senior) 0.95 (final agent) 4 HRL-2023 - 14 PPO / SAC 0.95 MARL-2023 - 5 MA-PPO / MA-SACD 0.9 HUGO-2024 - 118 PPO 0.85 (TopoAgent) 0.95 (final Senior) IL-2024 - 14 IL 0.97 HMO-2024 - 14, 118 A2C/PPO N.A. [^8] Not Available. It is either not documented or not used. N.A. N.A. BDQN-2024 - 14, 36 ApeX-DQN N.A. N.A. N.A. Zonal-2024 - 118 PPO 0.9 N.A. N.A. GNNIL-2025 - 14 IL 0.97 CCMA-2025 - 5, 14 PPO 0.95 SoftIL-2025 - 118 IL 0.9
@@ 3.1 General RL Techniques
3.1.1 Algorithm
As shown in Table 1, the most frequently used RL algorithm for PNC is PPO, featured in 11 of the 26 solutions reviewed in this overview, significantly more often than any other algorithm. Studies comparing PPO with other algorithms ([74, 47, 75]), such as SAC and A2C, consistently show that PPO performs better in PNC. Note, however, that the agent’s performance depends not only on the choice of algorithm but also on other implementation details, which are discussed in the following paragraphs. Additionally, various versions of DQN have been applied to the PNC problem. In [71] (D3QN-2022), the authors demonstrate that the prioritized D3QN, an advanced variant of DQN, significantly speeds up the training. To get an idea of the differences between these methods, we refer the reader to Section 1.3.2.
3.1.2 Action space
As outlined in Section 1, the actions available to operate the power network include bus splitting, line switching, generator re-dispatch or curtailment and, optionally, storage changes (introduced in [33]). Topological actions – bus splitting and line switching – serve as a low-cost option to alleviate thermal overloads on the network, as highlighted in [11]. However, the large, non-linear, and combinatorial nature of the action space makes optimal control of the grid topology beyond current state-of-the-art capabilities ([1, 34]), leaving these topological actions as a significant, yet underutilized, source of operational flexibility in power networks.
The inability of existing methods to navigate such a vast action space has motivated the L2RPN competitions (see [1, 11, 34]), where the primary goal is to use RL, to explore the full range of switching options to control power flows. As a result, topological actions are the main focus of agents used in L2RPN. Footnote 11 provides an overview of the number of feasible bus-splitting and line-switching actions per case. It shows how the number of bus-splitting actions increases drastically when going from case 14 to case 36.
@@ 3.1.2 Action space
also presented by [73]. Note that this computation also includes the actions of substations where only one configuration is possible due to connectivity constraints. Therefore, the result of this computation is slightly higher compared to what is presented in Footnote 11, which excludes these actions.
@@
Table 3: Action space reductions for case 14, the PyPowNet, realistic, and sandbox version. The first two solutions use case 14 in PyPowNet, which is not mentioned in Footnote 11.
In Tables 4 and 5 we omitted the “Original Size Action Space” column, as these solutions primarily focus on bus-splitting actions, $\mathcal{A}_{bus}$, for which the action space sizes are already detailed in Footnote 11. As can be seen in Table 4, the exceptions to this are A3C-2020 and D3QN-2020.
@@
Solution Acronym Actions for RL agent Reduced Size Action Space % of actions included Reduction method Actions for rule-based heuristics SMAAC-2021 Bus splitting 66674 99.8% Mask reduction with $N=5$. Line switching. A3C-2020 Topology actions 300 0.4% 200 based on experience, 100 randomly selected. In addition, the action space contains 295 line-switching actions. Line switching. SAS-2021 Bus splitting 732  [^15] Two action spaces are used, $\mathcal{A}_{normal}$, which is used by default and $\mathcal{A}_{exception}$, which is used only when specific lines are disconnected. Here, $|\mathcal{A}_{normal}|=500$ and $|\mathcal{A}_{exception}|=232$. 1.0% N.A. Line switching and redispatching. Binbinchen-2020 Bus splitting 208  [^16] This action space includes change bus actions. This is different from set bus actions which most solutions use. See https://grid2op.readthedocs.io/en/latest/. 0.3% Greedy reduction. Line switching when line is disconnected. D3QN-2020 Topology actions or redispatch actions 786 16  [^17] In addition, the action space contains 58 line switching, 40 redispatching and one explicit do-nothing action. 1.2% Expert reduction. Line switching when line is disconnected. PowRL-2022 Bus splitting 240 0.4% Greedy reduction. Line switching. Curriculum-2023 Bus splitting 508 16 0.8% Greedy reduction. Line switching. BDQN-2024 N.A., but only bus splitting actions seem to be used 1080 1.6% Stable Topo reduction. N.A.
@@
Solution Acronym Actions for RL agent Reduced size action space % of actions included Reduction method Actions for rule-based heuristics SAS-2021 Bus splitting 1000 1.0% N.A. Line switching. D3QN-2020 Topology actions 978 1.4% Expert reduction16. In addition, the action space contains 185 line switching and one do-nothing action. Line switching. AlphaZero-2022 Bus splitting 2000 2.7% Greedy reduction. Line switching and redispatching. BruteForce-2022 314 0.4% Greedy reduction. Bus splitting, line switching and redispatching. HRI-EU-2022 N.A. N.A. Randomly draw 1000 line switch combined with redispatch actions each time the grid is in danger. Line switching combined with redispatch. LJNAgent-2024 Bus splitting 1516 2.1% Greedy reduction. Line switching and redispatching. Artelys-2024 Bus splitting 500 0.7% Greedy reduction. Line switching and redispatching. HUGO-2024 Bus splitting 2030 2.8% Apply the action space of AlphaZero-2022 with an additional 30 expert actions selected by RTE (Expert reduction). Stable Topo Reduction for the target topologies (TTs). Line switching. SoftIL-2025 Bus splitting 2030 2.8% Apply the action space of HUGO-2024. Line switching.